OpenStudy (anonymous):

evaluate the integral of general odd function from negative infinity to infinity?

OpenStudy (anonymous):

\[I=\int\limits_{-\infty}^{\infty}f \left( x \right)~dx,where~f \left( x \right)~is~an~odd~function.\] put x=-y, dx=-dy \[when~x \rightarrow \infty,y \rightarrow-\infty ,when~x \rightarrow-\infty,y \rightarrow \infty \] \[I=\int\limits_{\infty}^{-\infty}f \left( -y \right)\left( -dy \right)=-\int\limits_{\infty}^{-\infty}f \left( -y \right)dy\] \[I=\int\limits_{-\infty}^{\infty}f \left( -y \right)dy=\int\limits_{-\infty}^{\infty}f \left( -x \right)dx=-\int\limits_{-\infty}^{\infty}f \left( x \right)dx=-I\] I+I=0 2I=0 I=0

OpenStudy (anonymous):

if f(x)is an odd function,then f(-x )=-f(x )